# Hurwitz moduli varieties parameterizing pointed covers of an algebraic curve with a fixed monodromy group

### v3

@@ 2. Parameterization of covers with a fixed monodromy group
The map $[f:X\to Y]\mapsto(D,[m]_{f})$ establishes a bijective correspondence between the set of equivalence classes of $(\Lambda,G)$-covers of $Y$ branched in $n$ points and the set $H_{n}^{\Lambda,G}(Y)$. If the centralizer of $G$ in $S(\Lambda)$ is trivial, $Z_{S(\Lambda)}(G)=\{1\}$, then the covering isomorphism between every two $(\Lambda,G)$-covers with the same monodromy invariant is unique.
@@
The following two conditions for a cover $f:X\to Y$ are equivalent
  1. (i)
@@
Choosing a marked element $\lambda_{0}\in\Lambda$ and imposing $\varepsilon(\lambda_{0})=x_{0}$ is a normalizing condition for pointed degree $d$ covers of $(Y,y_{0})$ whose monodromy group equals $G$. This serves for defining the monodromy invariant $(D,m^{N(\lambda_{0})})\in H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})$ so that to have the bijection of Proposition 2.13. The choice of another marked element $\lambda_{1}\in\Lambda$ is discussed in Proposition 4.23.
@@ 3. Smooth, proper families of covers with a fixed monodromy group
Under the assumptions of (i) the morphism $f:X\to Y\times S$ is finite, surjective and flat [21, Proposition 2.6]. If $S$ is connected, then all covers $f_{s}:X_{s}\to Y$ have the same degree $d$, where $d$ is the rank of the locally free sheaf $f_{\ast}\mathcal{O}_{X}$. It is clear that two families are equivalent if and only if there is an $S$-isomorphism $h:X_{1}\to X_{2}$, which is a covering isomorphism over $Y$ for $\forall s\in S$, i.e. $f_{1s}=f_{2s}\circ h_{s}$ for $\forall s\in S$.
@@
Let $D=\{b_{1},\ldots,b_{n}\}\subset Y\setminus y_{0}$, let $\overline{U}_{1},\ldots,\overline{U}_{n}$ be disjoint embedded closed disks in $Y\setminus y_{0}$ such that $b_{i}\in U_{i}$ for $\forall i$, where $U_{i}$ is the interior of $\overline{U}_{i}$. Let $N_{D}(U_{1},\ldots,U_{n})\subset(Y\setminus y_{0})^{(n)}_{\ast}$ be the open neighborhood of $D$ in the complex topology (i.e. that of $((Y\setminus y_{0})^{(n)}_{\ast})^{an}$), which consists of $E=\{y_{1},\ldots,y_{n}\}$ such that $y_{i}\in U_{i}$ for every $i$. The inclusion $Y\setminus\cup_{i=1}^{n}U_{i}\hookrightarrow Y\setminus D$ is a deformation retract, so for every homomorphism $m:\pi_{1}(Y\setminus D,y_{0})\to G$ and every $E\in N_{D}(U_{1},\ldots,U_{n})$ there is a unique homomorphism $m(E):\pi_{1}(Y\setminus E,y_{0})\to G$ such that the following diagram commutes
[…]
Given a path $\gamma:I\to Y\setminus E$ we denote by $[\gamma]_{E}$ its homotopy class in $Y\setminus E$. We denote by $N_{(D,m)}(U_{1},\ldots,U_{n})$ the subset of $H^{G}_{n}(Y,y_{0})$
(8) $N_{(D,m)}(U_{1},\ldots,U_{n})=\{(E,m(E))|E\in N_{D}(U_{1},\ldots,U_{n})\}$
@@
Let $D=\{b_{1},\ldots,b_{n}\}$ be the branch locus of $f_{s_{0}}$, let $y_{0}\in Y\setminus D$ and let $\varepsilon:\Lambda\to f_{s_{0}}^{-1}(y_{0})$, $m:\pi_{1}(Y\setminus D,y_{0})\twoheadrightarrow G$ satisfy $\varepsilon(\lambda m([\alpha]))=\varepsilon(\lambda)\alpha$ for $\forall\lambda\in\Lambda$ and $\forall[\alpha]\in\pi_{1}(Y\setminus D,y_{0})$. Let $B\subset Y\times S$ be the branch locus of $f$. The map $\beta:S\to Y^{(n)}_{\ast}$ defined by $\beta(s)=B_{s}$ is a morphism [21, Proposition 2.6], so $|\beta^{an}|:|S^{an}|\to|(Y^{(n)}_{\ast})^{an}|$ is a continuous map. Let $N_{D}(U_{1},\ldots,U_{n})$ be a neighborhood of $D=\beta(s_{0})$ as in § 3.4. Let $V_{1}$ be a neighborhood of $s_{0}$ in $|S^{an}|$ such that $\beta(V_{1})\subset N_{D}(U_{1},\ldots,U_{n})$. The restriction $|X^{an}|\setminus f^{-1}(B)\overset{|f^{an}|}{\longrightarrow}|(Y\times S)^{an}|\setminus B$ is a topological covering map [21, Proposition 2.6] and $(Y\times S)^{an}\cong Y^{an}\times S^{an}$ [27, § 1.2]. The complex space $S^{an}$ is locally connected (cf. [12, Ch.9 § 3 n.1]), therefore there is an embedded open disk $U\subset Y\setminus\cup_{i=1}^{n}\overline{U}_{i}$, $y_{0}\in U$ and a connected neighborhood $V$ of $s_{0}$, such that $V\subset V_{1}$, $U\times V\subset Y\times S\setminus B$ and $f^{-1}(U\times V)$ is a disjoint union of connected open sets homeomorphic to $U\times V$. Denote these open sets by $W_{\lambda}$, $\lambda\in\Lambda$, where $W_{\lambda}\ni\varepsilon(\lambda)$. For every $s\in V$ we define a bijection $\varepsilon_{s}:\Lambda\to f^{-1}(y_{0},s)$ and an epimorphism $m_{s}:\pi_{1}(Y\setminus B_{s},y_{0})\to G$ by
@@
[…]
@@
$\varepsilon_{s}(\lambda m_{s}([\alpha]_{B_{s}}))=\varepsilon_{s}(\lambda)\alpha\quad\text{for}\>\forall\lambda\in\Lambda\;\text{and}\;\forall[\alpha]\in\pi_{1}(Y\setminus B_{s},y_{0})$
@@
Let $d=|\Lambda|$. Every cover $f_{s}:X_{s}\to Y$, $s\in S$, is of degree $d$ since $S$ is connected. We have to prove that $f_{s}:X_{s}\to Y$ is a $(\Lambda,G)$-cover for $\forall s\in S$. For every transitive subgroup $H\subset S(\Lambda)$ let $S^{H}$ be the set of points $s\in S$ such that $f_{s}:X_{s}\to Y$ is a $(\Lambda,H)$-cover. It is clear that $S=\cup_{H}S^{H}$. By Lemma 3.5 $S^{H}$ is an open set in $|S^{an}|$ for every $H$. By Proposition 2.20, given two transitive subgroups $H$ and $K$ of $S(\Lambda)$, the following alternative holds: if $H$ and $K$ are conjugated in $S(\Lambda)$, then $S^{H}=S^{K}$; otherwise $S^{H}\cap S^{K}=\emptyset$. The topological space $|S^{an}|$ is connected (cf. [27, Corollaire 2.6]), therefore $S=S^{G}$. ∎
@@
Let $S$ be an algebraic variety. We denote by $\mathcal{H}^{\Lambda,G}_{Y,n}(S)$ the set of equivalence classes $[f:X\to Y\times S]$ of smooth, proper families of $(\Lambda,G)$-covers of $Y$ branched in $n$ points, parameterized by $S$ (cf. Definition 3.1). We denote by ($\mathcal{H}^{\Lambda,G}_{(Y,y_{0}),n}(S)$ the set of equivalence classes $[f:X\to Y\times S,\eta:S\to X]$ of smooth, proper families of pointed $(\Lambda,G)$-covers of $(Y,y_{0})$ branched in $n$ points parameterized by $S$ (cf. Definition 3.2)
@@
Let $u:T\to S$ be a morphism of algebraic varieties. Given a smooth, proper morphism $X\to S$ of reduced, separated schemes of finite type over $\mathbb{C}$, the pullback morphism $X_{T}=X\times_{S}T\to T$ is smooth and proper. The scheme $X\times_{S}T$ is reduced since $T$ is reduced (cf. [24, p.184]. Hence $X\times_{S}T$ is isomorphic to the closed algebraic subvariety of $X\times T$ whose set of points is the set-theoretical fiber product $X(\mathbb{C})\times_{S(\mathbb{C})}T(\mathbb{C})$.
Let $f:X\to Y\times S$ be a smooth, proper family of covers of $Y$ branched in $n$ points. Let $u:T\to S$ be a morphism of algebraic varieties. Let $X_{T}\to T$ be the pullback of $\pi_{2}\circ f:X\to S$. By the above argument $X_{T}=\{(x,t)|\pi_{2}\circ f(x)=u(t)\}\subset X\times T$. Let $f_{T}:X_{T}\to Y\times T$ and $h:X_{T}\to X$ be the morphisms defined respectively by $f_{T}(x,t)=(\pi_{1}\circ f(x),t)$ and $h(x,t)=x$. One has the following commutative diagram of morphisms
[…]
@@
Given a smooth, proper family $(f:X\to Y\times S,\eta:S\to X)$ of pointed $(\Lambda,G)$-covers of $(Y,y_{0})$ branched in $n$ points and a morphism $u:T\to S$ one defines the pullback family as $(f_{T}:X_{T}\to Y\times T,\eta_{T}:T\to X_{T})$, where $\eta_{T}$ is the morphism defined by $\eta_{T}(t)=(\eta(u(t)),t)$. This is a smooth, proper family of pointed $(\Lambda,G)$-covers of $(Y,y_{0})$. This defines, as above, a moduli functor ($\mathcal{H}^{\Lambda,G}_{(Y,y_{0}),n}:Var_{\mathbb{C}}\to(Sets)$.
@@
Let $y_{0}\in Y$. If furthermore $p_{s}:\mathcal{C}_{s}\to Y\times\{s\}$ is unbranched at $(y_{0},s)$ for $\forall s\in S$ and there exists a morphism $\zeta:S\to\mathcal{C}$ such that $p\circ\zeta(s)=(y_{0},s)$ for $\forall s\in S$, then $(p:\mathcal{C}\to Y\times S,\zeta)$ is called a smooth, proper family of pointed $G$-covers of $(Y,y_{0})$ branched in $n$ points (cf. [21, Section 3]).
@@
The group $G$ acts by covering automorphisms of the finite morphism $\Lambda\times\mathcal{C}\to Y\times S$ defined by $(\lambda,z)\mapsto p(z)$. Therefore $X=(\Lambda\times\mathcal{C})/G$ has a structure of quotient algebraic variety and the maps $\pi:\Lambda\times\mathcal{C}\to X$ and $f:X\to Y\times S$ are finite morphisms [28, Ch.III Prop. 19]. The morphism $\pi_{2}\circ f:X\to S$ is a composition of a proper and a finite morphism, so it is proper. For every $s\in S$ its scheme-theoretical fiber $X_{s}$ is isomorphic to the quotient $(\Lambda\times\mathcal{C}_{s})/G$, since formation of quotients by $G$ commutes with base change (cf. [22, Prop. A.7.1.3]). Therefore $X_{s}$ is smooth and irreducible. Let $H=G(\lambda_{0})$. Similar statements hold for $X_{1}=\mathcal{C}/H$, $\rho:\mathcal{C}\to X_{1}$ and $f_{1}:X_{1}\to Y\times S$. The two families of covers of $Y$, $f:X\to Y\times S$ and $f_{1}:X_{1}\to Y\times S$ are equivalent. In fact the morphisms $\mathcal{C}\to\Lambda\times^{G}\mathcal{C}=X$ and $\Lambda\times\mathcal{C}\to\mathcal{C}/H=X_{1}$, defined as in (5), induce morphisms $X_{1}\to X$ and $X\to X_{1}$ over $Y\times S$, given respectively by $\rho(z)\mapsto\pi(\lambda_{0},z)$ and $\pi(\lambda,z)=\rho(a_{\lambda}z)$, which are inverse to each other. We claim that $\pi_{2}\circ f$ and $\pi_{2}\circ f_{1}$ are flat morphisms. Let $x\in X_{1}$, let $\pi_{2}\circ f_{1}(x)=s$ and let $\rho^{-1}(x)$ be the scheme-theoretical fiber of $\rho$. The restriction $\rho_{s}:\mathcal{C}_{s}\to(X_{1})_{s}$ is a surjective morphism of smooth, irreducible, projective curves, so $\dim_{\mathbb{C}}H^{0}(\mathcal{O}_{\rho_{s}^{-1}(x)})=|H|=|G|/|\Lambda|$. The schemes $\rho^{-1}(x)$ and $\rho_{s}^{-1}(x)$ coincide, so $\dim_{\mathbb{C}}H^{0}(\mathcal{O}_{\rho^{-1}(x)})=|G|/|\Lambda|$ for every $x\in X_{1}$. This implies that the coherent sheaf $\rho_{\ast}\mathcal{O}_{\mathcal{C}}$ is locally free (cf. [26, Ch. 2, § 5, Lemma 1]), therefore $\rho:\mathcal{C}\to X_{1}$ is flat. It is moreover faithfully flat since $\rho(\mathcal{C})=X_{1}$. By hypothesis $\pi_{2}\circ p:\mathcal{C}\to S$ is flat, therefore $\pi_{2}\circ f_{1}:X_{1}\to S$ is flat (cf. [24, p.46]). As we saw above the scheme-theoretical fiber $(X_{1})_{s}$ over every (closed) point $s\in S$ is smooth. Therefore $\pi_{2}\circ f_{1}:X_{1}\to S$ is smooth at every closed point of $X_{1}$ (cf. [2, Ch.VII, Thm.(1.8)]. The points of the scheme $X_{1}$ at which $\pi_{2}\circ f_{1}$ is smooth form an open subset. This open subset contains every closed point of $X_{1}$, therefore it coincides with $X_{1}$ (cf. [11, Prop. 3.35]). The smoothness of $\pi_{2}\circ f_{1}:X_{1}\to S$ implies the smoothness of $\pi_{2}\circ f:X\to S$ since there is an $Y\times S$-isomorphism between $X$ and $X_{1}$. For every $s\in S$ one has $X_{s}\cong(\Lambda\times\mathcal{C}_{s})/G$, so the morphism $f_{s}:X_{s}\to Y$, is a $(\Lambda,G)$-cover of $Y$ whose branch locus coincides with that of $p_{s}:\mathcal{C}_{s}\to Y$ (cf. Lemma 2.12). Therefore $f:X\to Y\times S$ is a smooth, proper family of $(\Lambda,G)$-covers of $Y$ branched in $n$ points and it is equivalent to $f_{1}:X_{1}\to Y\times S$ as we saw above. ∎

### v4

@@ 2. Parameterization of covers with a fixed monodromy group
The map $f:X\to Y]\mapsto(D,[m]_{f})$ establishes a bijective correspondence between the set of equivalence classes of $(\Lambda,G)$-covers of $Y$ branched in $n$ points and the set $H_{n}^{\Lambda,G}(Y)$. If the centralizer of $G$ in $S(\Lambda)$ is trivial, $Z_{S(\Lambda)}(G)=\{1\}$, then the covering isomorphism between every two $(\Lambda,G)$-covers with the same monodromy invariant is unique.
@@
The following two conditions for a cover $f:X\to Y$ are equivalent:
  1. (1)
@@
Choosing a marked element $\lambda_{0}\in\Lambda$ and imposing $\varepsilon(\lambda_{0})=x_{0}$ is a normalizing condition for pointed degree $d$ covers of $(Y,y_{0})$ whose monodromy group equals $G$. This serves for defining the monodromy invariant $(D,m^{N(\lambda_{0})})\in H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})$ so that the bijection of Proposition 2.13 to hold. The choice of another marked element $\lambda_{1}\in\Lambda$ is discussed in Proposition 4.23.
@@ 3. Smooth, proper families of covers with a fixed monodromy group
Under the assumptions of (i) the morphism $f:X\to Y\times S$ is finite, surjective and flat [22, Proposition 2.6]. If $S$ is connected, then all covers $f_{s}:X_{s}\to Y$ have the same degree $d$, where $d$ is the rank of the locally free sheaf $f_{\ast}\mathcal{O}_{X}$. It is clear that two families are equivalent if and only if there is an $S$-isomorphism $h:X_{1}\to X_{2}$, which is a covering isomorphism over $Y$ for $\forall s\in S$, i.e. $f_{1s}=f_{2s}\circ h_{s}$ for $\forall s\in S$.
@@
Let $D=\{b_{1},\ldots,b_{n}\}\subset Y\setminus y_{0}$, let $\overline{U}_{1},\ldots,\overline{U}_{n}$ be disjoint embedded closed disks in $Y\setminus y_{0}$ such that $b_{i}\in U_{i}$ for $\forall i$, where $U_{i}$ is the interior of $\overline{U}_{i}$. Let $N_{D}(U_{1},\ldots,U_{n})\subset(Y\setminus y_{0})^{(n)}_{\ast}$ be the open neighborhood of $D$ in the complex topology (i.e. that of $((Y\setminus y_{0})^{(n)}_{\ast})^{an}$), which consists of $E=\{y_{1},\ldots,y_{n}\}$ such that $y_{i}\in U_{i}$ for every $i$. The inclusion $Y\setminus\cup_{i=1}^{n}U_{i}\hookrightarrow Y\setminus D$ is a deformation retract, so for every homomorphism $m:\pi_{1}(Y\setminus D,y_{0})\to G$ and every $E\in N_{D}(U_{1},\ldots,U_{n})$ there is a unique homomorphism $m(E):\pi_{1}(Y\setminus E,y_{0})\to G$ such that the following diagram commutes:
${\pi_{1}(Y\setminus D,y_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{m}$${\pi_{1}(Y\setminus\cup_{i=1}^{n}U_{i},y_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}\scriptstyle{\cong}$${\pi_{1}(Y\setminus E,y_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{m(E)}$${G}$
Given a path $\gamma:I\to Y\setminus E$ we denote by $\gamma]_{E}$ its homotopy class in $Y\setminus E$. We denote by $N_{(D,m)}(U_{1},\ldots,U_{n})$ the subset of $H^{G}_{n}(Y,y_{0})$
(8) $N_{(D,m)}(U_{1},\ldots,U_{n})=\{(E,m(E))|E\in N_{D}(U_{1},\ldots,U_{n})\}.$
@@
Let $D=\{b_{1},\ldots,b_{n}\}$ be the branch locus of $f_{s_{0}}$, let $y_{0}\in Y\setminus D$ and let $\varepsilon:\Lambda\to f_{s_{0}}^{-1}(y_{0})$, $m:\pi_{1}(Y\setminus D,y_{0})\twoheadrightarrow G$ satisfy $\varepsilon(\lambda m([\alpha]))=\varepsilon(\lambda)\alpha$ for $\forall\lambda\in\Lambda$ and $\forall[\alpha]\in\pi_{1}(Y\setminus D,y_{0})$. Let $B\subset Y\times S$ be the branch locus of $f$. The map $\beta:S\to Y^{(n)}_{\ast}$ defined by $\beta(s)=B_{s}$ is a morphism [22, Proposition 2.6], so $|\beta^{an}|:|S^{an}|\to|(Y^{(n)}_{\ast})^{an}|$ is a continuous map. Let $N_{D}(U_{1},\ldots,U_{n})$ be a neighborhood of $D=\beta(s_{0})$ as in §3.4. Let $V_{1}$ be a neighborhood of $s_{0}$ in $|S^{an}|$ such that $\beta(V_{1})\subset N_{D}(U_{1},\ldots,U_{n})$. The restriction $|X^{an}|\setminus f^{-1}(B)\overset{|f^{an}|}{\longrightarrow}|(Y\times S)^{an}|\setminus B$ is a topological covering map [22, Proposition 2.6] and $(Y\times S)^{an}\cong Y^{an}\times S^{an}$ [28, §1.2]. The complex space $S^{an}$ is locally connected (cf. [12, Ch.9 §3 n.1]), therefore there is an embedded open disk $U\subset Y\setminus\cup_{i=1}^{n}\overline{U}_{i}$, $y_{0}\in U$ and a connected neighborhood $V$ of $s_{0}$, such that $V\subset V_{1}$, $U\times V\subset Y\times S\setminus B$ and $f^{-1}(U\times V)$ is a disjoint union of connected open sets homeomorphic to $U\times V$. Denote these open sets by $W_{\lambda}$, $\lambda\in\Lambda$, where $W_{\lambda}\ni\varepsilon(\lambda)$. For every $s\in V$ we define a bijection $\varepsilon_{s}:\Lambda\to f^{-1}(y_{0},s)$ and an epimorphism $m_{s}:\pi_{1}(Y\setminus B_{s},y_{0})\to G$ by
@@
(cf. §3.4). We claim that $(\varepsilon_{s},m_{s})$ satisfies Condition (i) of Proposition 2.15 for $f_{s}:X_{s}\to Y$ (we identify $Y\times\{s\}$ with $Y$). Let $\alpha$ be a loop in $Y\setminus\cup_{i=1}^{n}U_{i}$ based at $y_{0}$. Let $g=m([\alpha]_{D})$. We claim that $\varepsilon_{s}(\lambda)\alpha=\varepsilon_{s}(\lambda g)$ for $\forall\lambda\in\Lambda$ and $\forall s\in V$. Consider the homotopy $F:[0,1]\times V\to Y^{an}\times S^{an}\setminus B$ defined by $F(t,s)=(\alpha(t),s)$. Let $\lambda\in\Lambda$. The restriction $f|_{W_{\lambda}}:W_{\lambda}\to U\times V$ is a homeomorphism. Let us denote by $\tilde{F}_{\lambda}(0):\{0\}\times V\to|X^{an}\setminus f^{-1}(B)|$ the composition $(0,s)\overset{F}{\mapsto}(y_{0},s)\mapsto(f|_{W_{\lambda}})^{-1}(y_{0},s)=\varepsilon_{s}(\lambda)$. By the covering homotopy property (cf. [30, Ch.2 §3 Th.3]) there is a unique continuous lifting $\tilde{F}_{\lambda}$ of $F$ which extends $\tilde{F}_{\lambda}(0)$
${X^{an}\setminus f^{-1}(B)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${[0,1]\times V\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\tilde{F}_{\lambda}}$$\scriptstyle{F}$${Y^{an}\times S^{an}\setminus B.}$
@@
$\varepsilon_{s}(\lambda m_{s}([\alpha]_{B_{s}}))=\varepsilon_{s}(\lambda)\alpha\quad\text{for}\>\forall\lambda\in\Lambda\;\text{and}\;\forall[\alpha]\in\pi_{1}(Y\setminus B_{s},y_{0}).$
@@
Let $d=|\Lambda|$. Every cover $f_{s}:X_{s}\to Y$, $s\in S$, is of degree $d$ since $S$ is connected. We have to prove that $f_{s}:X_{s}\to Y$ is a $(\Lambda,G)$-cover for $\forall s\in S$. For every transitive subgroup $H\subset S(\Lambda)$ let $S^{H}$ be the set of points $s\in S$ such that $f_{s}:X_{s}\to Y$ is a $(\Lambda,H)$-cover. It is clear that $S=\cup_{H}S^{H}$. By Lemma 3.5 $S^{H}$ is an open set in $|S^{an}|$ for every $H$. By Proposition 2.20, given two transitive subgroups $H$ and $K$ of $S(\Lambda)$, the following alternative holds: if $H$ and $K$ are conjugated in $S(\Lambda)$, then $S^{H}=S^{K}$; otherwise $S^{H}\cap S^{K}=\emptyset$. The topological space $|S^{an}|$ is connected (cf. [28, Corollaire 2.6]), therefore $S=S^{G}$. ∎
@@
Let $S$ be an algebraic variety. We denote by $\mathcal{H}^{\Lambda,G}_{Y,n}(S)$ the set of equivalence classes $f:X\to Y\times S]$ of smooth, proper families of $(\Lambda,G)$-covers of $Y$ branched in $n$ points, parameterized by $S$ (cf. Definition 3.1). We denote by $\mathcal{H}^{\Lambda,G}_{(Y,y_{0}),n}(S)$ the set of equivalence classes $f:X\to Y\times S,\eta:S\to X]$ of smooth, proper families of pointed $(\Lambda,G)$-covers of $(Y,y_{0})$ branched in $n$ points parameterized by $S$ (cf. Definition 3.2)
@@
Let $u:T\to S$ be a morphism of algebraic varieties. Given a smooth, proper morphism $X\to S$ of reduced, separated schemes of finite type over $\mathbb{C}$, the pullback morphism $X_{T}=X\times_{S}T\to T$ is smooth and proper. The scheme $X\times_{S}T$ is reduced since $T$ is reduced (cf. [25, p.184]. Hence $X\times_{S}T$ is isomorphic to the closed algebraic subvariety of $X\times T$ whose set of points is the set-theoretical fiber product $X(\mathbb{C})\times_{S(\mathbb{C})}T(\mathbb{C})$.
Let $f:X\to Y\times S$ be a smooth, proper family of covers of $Y$ branched in $n$ points. Let $u:T\to S$ be a morphism of algebraic varieties. Let $X_{T}\to T$ be the pullback of $\pi_{2}\circ f:X\to S$. By the above argument $X_{T}=\{(x,t)|\pi_{2}\circ f(x)=u(t)\}\subset X\times T$. Let $f_{T}:X_{T}\to Y\times T$ and $h:X_{T}\to X$ be the morphisms defined respectively by $f_{T}(x,t)=(\pi_{1}\circ f(x),t)$ and $h(x,t)=x$. One has the following commutative diagram of morphisms:
(11) ${X_{T}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{T}}$$\scriptstyle{h}$${Y\times T\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{id\times u}$$\scriptstyle{p_{2}}$${T\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{u}$${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${Y\times S\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{2}}$${S}$
@@
Given a smooth, proper family $(f:X\to Y\times S,\eta:S\to X)$ of pointed $(\Lambda,G)$-covers of $(Y,y_{0})$ branched in $n$ points and a morphism $u:T\to S$ one defines the pullback family as $(f_{T}:X_{T}\to Y\times T,\eta_{T}:T\to X_{T})$, where $\eta_{T}$ is the morphism defined by $\eta_{T}(t)=(\eta(u(t)),t)$. This is a smooth, proper family of pointed $(\Lambda,G)$-covers of $(Y,y_{0})$. This defines, as above, a moduli functor $\mathcal{H}^{\Lambda,G}_{(Y,y_{0}),n}:Var_{\mathbb{C}}\to(Sets)$.
@@
Let $y_{0}\in Y$. If furthermore $p_{s}:\mathcal{C}_{s}\to Y\times\{s\}$ is unbranched at $(y_{0},s)$ for $\forall s\in S$ and there exists a morphism $\zeta:S\to\mathcal{C}$ such that $p\circ\zeta(s)=(y_{0},s)$ for $\forall s\in S$, then $(p:\mathcal{C}\to Y\times S,\zeta)$ is called a smooth, proper family of pointed $G$-covers of $(Y,y_{0})$ branched in $n$ points (cf. [22, Section 3]).
@@
The group $G$ acts by covering automorphisms of the finite morphism $\Lambda\times\mathcal{C}\to Y\times S$ defined by $(\lambda,z)\mapsto p(z)$. Therefore $X=(\Lambda\times\mathcal{C})/G$ has a structure of quotient algebraic variety and the maps $\pi:\Lambda\times\mathcal{C}\to X$ and $f:X\to Y\times S$ are finite morphisms [29, Ch.III Prop. 19]. The morphism $\pi_{2}\circ f:X\to S$ is a composition of a proper and a finite morphism, so it is proper. For every $s\in S$ its scheme-theoretical fiber $X_{s}$ is isomorphic to the quotient $(\Lambda\times\mathcal{C}_{s})/G$, since formation of quotients by $G$ commutes with base change (cf. [23, Prop. A.7.1.3]). Therefore $X_{s}$ is smooth and irreducible. Let $H=G(\lambda_{0})$. Similar statements hold for $X_{1}=\mathcal{C}/H$, $\rho:\mathcal{C}\to X_{1}$ and $f_{1}:X_{1}\to Y\times S$. The two families of covers of $Y$, $f:X\to Y\times S$ and $f_{1}:X_{1}\to Y\times S$ are equivalent. In fact the morphisms $\mathcal{C}\to\Lambda\times^{G}\mathcal{C}=X$ and $\Lambda\times\mathcal{C}\to\mathcal{C}/H=X_{1}$, defined as in (5), induce morphisms $X_{1}\to X$ and $X\to X_{1}$ over $Y\times S$, given respectively by $\rho(z)\mapsto\pi(\lambda_{0},z)$ and $\pi(\lambda,z)=\rho(a_{\lambda}z)$, which are inverse to each other. We claim that $\pi_{2}\circ f$ and $\pi_{2}\circ f_{1}$ are flat morphisms. Let $x\in X_{1}$, let $\pi_{2}\circ f_{1}(x)=s$ and let $\rho^{-1}(x)$ be the scheme-theoretical fiber of $\rho$. The restriction $\rho_{s}:\mathcal{C}_{s}\to(X_{1})_{s}$ is a surjective morphism of smooth, irreducible, projective curves, so $\dim_{\mathbb{C}}H^{0}(\mathcal{O}_{\rho_{s}^{-1}(x)})=|H|=|G|/|\Lambda|$. The schemes $\rho^{-1}(x)$ and $\rho_{s}^{-1}(x)$ coincide, so $\dim_{\mathbb{C}}H^{0}(\mathcal{O}_{\rho^{-1}(x)})=|G|/|\Lambda|$ for every $x\in X_{1}$. This implies that the coherent sheaf $\rho_{\ast}\mathcal{O}_{\mathcal{C}}$ is locally free (cf. [27, Ch. 2, §5, Lemma 1]), therefore $\rho:\mathcal{C}\to X_{1}$ is flat. It is moreover faithfully flat since $\rho(\mathcal{C})=X_{1}$. By hypothesis $\pi_{2}\circ p:\mathcal{C}\to S$ is flat, therefore $\pi_{2}\circ f_{1}:X_{1}\to S$ is flat (cf. [25, p.46]). As we saw above the scheme-theoretical fiber $(X_{1})_{s}$ over every (closed) point $s\in S$ is smooth. Therefore $\pi_{2}\circ f_{1}:X_{1}\to S$ is smooth at every closed point of $X_{1}$ (cf. [2, Ch.VII, Thm.(1.8)]. The points of the scheme $X_{1}$ at which $\pi_{2}\circ f_{1}$ is smooth form an open subset. This open subset contains every closed point of $X_{1}$, therefore it coincides with $X_{1}$ (cf. [11, Prop. 3.35]). The smoothness of $\pi_{2}\circ f_{1}:X_{1}\to S$ implies the smoothness of $\pi_{2}\circ f:X\to S$ since there is an $Y\times S$-isomorphism between $X$ and $X_{1}$. For every $s\in S$ one has $X_{s}\cong(\Lambda\times\mathcal{C}_{s})/G$, so the morphism $f_{s}:X_{s}\to Y$, is a $(\Lambda,G)$-cover of $Y$ whose branch locus coincides with that of $p_{s}:\mathcal{C}_{s}\to Y$ (cf. Lemma 2.12). Therefore $f:X\to Y\times S$ is a smooth, proper family of $(\Lambda,G)$-covers of $Y$ branched in $n$ points and it is equivalent to $f_{1}:X_{1}\to Y\times S$ as we saw above. ∎